OpenStudy (bekkah323):

120=rt 120=2r(t-30) Solve for R

OpenStudy (anonymous):

use substitution method

OpenStudy (anonymous):

r=120/t from first equation solving for r 120=2(120/t )(t-30) plug in r into second equation 120=2(120/t )(t-30) now solve for t t= 60 plug that back into the top equation 120=60*r 120/60=r r= 2 ALSO i think it would have been faster if you solve for t first then plug in to equation 2 then solve for r and ended there :D